# Understanding and Mitigating Miscalibration in Prompt Tuning for Vision-Language Models

### v1

@@ 4 Method: Dynamic Outlier Regularization
In the previous analysis, we show that the textual divergence is the key to the calibration performance of fine-tuned CLIP. To preserve the calibration of zero-shot CLIP, our key idea is to regularize the textual divergence of new classes without restricting those of base classes. To this end, we propose to utilize textual outliers to improve the reliability of fine-tuned CLIP. Such a regularization can mitigate the overconfidence in the new classes, which are not explicitly inclusive in the fine-tuning.
@@
With this in mind, we construct a set of text outliers using nouns from WordNet (Miller, 1995), a large English lexical database containing over 150,000 words. We select nouns from WordNet that do not overlap but share higher-level concept relations with the base classes used in the fine-tuning, and then incorporate them into our regularization term for prompt tuning. We demonstrate the effectiveness of using relevant text outliers in Table 5.
@@
$s_{i}=\frac{1}{n}\sum_{j=1}^{n}\operatorname{sim}\left(\psi\left(\bm{t}_{o_{i}}\right),\psi\left(\bm{t}_{c_{j}}\right)\right),\text{ where }i\in\{1,2,\ldots,m\},$ (6)
where $\bm{t}_{o_{i}}$ represents the textual tokens of a noun $o_{i}$ using a fixed prompt like “a photo of a [class-name]”, and $\psi$ is the text encoder of zero-shot CLIP. Then we get the set of textual outliers using the score ranking.
@@
$\mathcal{L}_{\text{dor}}=1-\frac{1}{B}\sum_{b=1}^{B}\operatorname{sim}\left(\psi\left(\bm{t}_{o_{b}}^{\prime}\right),\psi\left(\bm{t}_{o_{b}}\right)\right).$ (8)
where $\operatorname{sim(\cdot)}$ denotes the cosine similarity function and $\bm{t}_{o_{b}}$ denotes the token of the textual outlier $o_{b}$. Using the regularization, the textual divergence of the fine-tuned CLIP will be regularized to be consistent with the zero-shot CLIP. Different from KgCoOp Yao et al. (2023), the outlier regularization does not restrict the textual feature deviation of base classes, which is explicitly shown in Figure 5.
@@
$\mathcal{L}_{total}=\mathcal{L}_{\text{ce}}+\lambda\cdot\mathcal{L}_{\text{dor}}.$ (9)
@@
Extension to other fine-tuning algorithms
It is worth noting that our regularization is a general method and can be easily incorporated into existing prompt tuning algorithms for CLIP, including knowledge-guided fine-tuning (Yao et al., 2023; 2024), multimodal consistency (Khattak et al., 2023), Decoupled prompt tuning (Zhang et al., 2024) etc. Given the vanilla fine-tuning objective $\mathcal{L}_{\text{base}}$ of the other methods and the hyperparameter $\lambda$, we formalize the fine-tuning objective as:
$\mathcal{L}_{\text{total }}=\mathcal{L}_{\text{base}}+\lambda\cdot\mathcal{L}_{\text{dor}}.$ (10)
Noticeably, DOR derived from textual outliers offers several compelling advantages:
  • Algorithm-agnostic: DOR can be easily incorporated into existing prompt tuning methods and consistently mitigate the calibration error under various evaluations (See Table 1 and 2). Furthermore, our method can be extended to visual fine-tuning methods with image outliers (See Section 4).
  • Easy-to-use: DOR leverages text outlier as the data for regularization, which is readily available and easy to collect. In Section 4 we will verify that DOR outperforms other selection strategies. We provide the detailed selection result of text outliers in Appendix C.
@@ 5 Experiments
5.1 Experimental Setup
Benchmark setting.
Following recent works (Zhou et al., 2022b; Wang et al., 2024), we perform two evaluations in two standard benchmark settings: 1) Generalization from Base-to-New Classes: A downstream dataset will be equally split into base and new classes. The model is trained only on the base classes in a few-shot setting and evaluated on base and new categories. 2) Domain Generalization: The modal is trained on ImageNet-1K in a few-shot manner and evaluated on four other ImageNet datasets that contain various types of domain shifts.
Datasets.
For the base-to-new evaluation, we use 11 image recognition datasets that cover diverse classification tasks including ImageNet (Deng et al., 2009), Caltech101 (Fei-Fei et al., 2004), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), FGVCAircraft (Maji et al., 2013), SUN397 (Xiao et al., 2010), UCF101 (Soomro et al., 2012), DTD (Cimpoi et al., 2014) and EuroSAT (Helber et al., 2019). For domain generalization, we use ImageNet-1K as the source dataset and its four variants as target datasets including ImageNetV2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a).
Baselines.
We compare and incorporate our method with existing fine-tuning algorithms including CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), KgCoOp (Yao et al., 2023), MaPLe (Khattak et al., 2023), DEPT (Zhang et al., 2024), and TCP (Yao et al., 2024).
Implementation details.
We use CLIP ( ViT-B/16) (Radford et al., 2021) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022a). For the compared tuning methods, we adopt them from the corresponding official implementation. For hyperparameter $\lambda$ in DOR, we set $\lambda$ = 8.0 for CoOp and 2.0 for other fine-tuning methods. We set the number of selected dynamic outlier repository to 5000. The number of outliers in each batch is the same as the base classes. We list the details of the compared methods in Appendix B.
Evaluation metrics.
Four standard metrics of confidence calibration are used in our evaluation, including Expected Calibration Error (ECE) (Guo et al., 2017), Maximum Calibration Error (MCE) (Guo et al., 2017), Adaptive Calibration Error (ACE) (Nixon et al., 2019) and Proximity-Informed Expected Calibration Error (PIECE) (Xiong et al., 2023).
Table 1: Average calibration performance across 11 datasets. “+DOR(Ours)” to our method applied to existing tuning methods. ↓ indicates smaller values are better. Calibration error is given by ×$\times 10^{-2}$. “HM” denotes the harmonic mean. Bold numbers are significantly superior results.
@@ 5.1 Experimental Setup
TCP 4.71 4.07 4.39 4.73 4.03 4.38 1.28 1.21 1.24 6.06 6.29 6.18
+DOR(Ours) 4.79 3.80 4.29 4.76 3.71 4.24 1.28 1.11 1.20 6.00 6.22 6.11
@@ 5.2 Results
Generally, we investigate the effectiveness of DOR from three metrics: calibration, generalization accuracy, and confidence level. Due to space constraints, we provide detailed calibration results of all datasets in Appendix D.
Table 2: Average base-to-new accuracy (%) of six methods with DOR on 11 datasets. “Vanilla” denotes the baseline of fine-tuning methods. Bold numbers are significantly superior results.
ZSCLIP CoOp CoCoOp MaPLe KgCoOp DEPT TCP
Base 69.49 82.97 83.20 80.57 79.89 82.11 82.08 82.29 82.13 83.70 83.81 83.95 83.89
New 74.32 61.74 72.01 72.47 74.59 73.89 75.89 72.21 73.14 65.04 71.39 75.21 75.31
HM 71.90 72.36 77.61 76.52 77.24 78.00 78.98 77.25 77.64 74.37 77.60 79.58 79.60
DOR enhances the overall calibration of existing prompt-tuning methods.
@@
Table 1 shows the calibration performance of the six baselines w/ or w/o our DOR framework under the base-to-new evaluation protocol. From the average results in the table, we observe a trade-off between the base and new class for all of the baseline methods, e.g., CoOp outperforms TCP on base classes but significantly underperforms TCP on new classes. Notably, DOR consistently reduces miscalibration on new classes across various metrics without calibration trade-offs on base and new classes. For instance, DOR significantly reduces the ECE from 14.58% to 6.49% on new classes and maintains the ECE from 3.07% to 2.67% on base classes, which makes CoOp more reliable in terms of predicted confidence. To showcase the versatility of DOR, we also present that DOR is also applicable to multi-modal tuning (e.g., MaPLe). MaPLe adjusts both vision and language branches and DOR consistently improves the calibration performance on new classes of it. In summary, our proposed DOR can consistently boost calibration performance on new classes upon existing state-of-the-art prompt tuning methods without compromising the vanilla fine-tuning objectives.
DOR benefits base-to-new generalization. To further verify that our DOR is effective on new classes and non-toxic for performance on base classes, we summarize the comparison of average test accuracy in Table 2. Similar to the evaluation of calibration, a salient observation is that our proposed DOR drastically improves base-to-new generalization, with its accuracy consistently outperforming all existing baselines in the harmonic mean of base and new classes. Moreover, DOR almost entirely preserves model capability in the classification performance on base classes without degeneration. For instance, applying DOR in DEPT can increase accuracy on new classes from 65.04% to 71.39%, while keeping the accuracy of base classes similar to the baseline. Given that most prompt tuning methods lag behind zero-shot CLIP on the accuracy of new classes, an intuitive explanation is that DOR aligns the features of unseen classes with the zero-shot features, which can preserve the zero-shot generalization on new clasess. In addition, we observe that some methods (e.g., CoOp and MaPLe) with DOR, resulting in improvements of 2.12% and 2.00% respectively, outperforming zero-shot accuracy on new classes. This demonstrates that DOR can significantly enhance the base-to-new generalization capacity of fine-tuned CLIP.
Figure 4: Comparison between KgCoOp (KG) and DOR (W/ CoOp). Left: Accuracy. Right: ECE.
DOR modifies the logit distribution and confidence level. To further illustrate the influence of DOR on confidence calibration, we visualize and compare the distribution of output logit and softmax confidence score for base and new classes in Figure 5. We compare zero-shot CLIP and CoOp w/ or w/o DOR on FGVCAircraft dataset. We can observe that if the model tuning With CoOp+DOR, the logit distribution of base class approximate CoOp, and the new class is similar to zero-shot CLIP, respectively. A similar phenomenon is observed in the softmax probability distribution. The results meet our vision mentioned in Section 3.1, DOR can leverage the advantages of both models, which ensure the confidence calibration on both base and new classes after fine-tuning.
@@
Table 3: Comparison of prompt learning in the domain generalization of ImageNet. DOR boosts the performance of existing methods on calibration and generalization.
Accuracy (↑) ECE (↓)
@@
CLIP 66.73 60.87 46.09 47.81 73.98 57.19 1.86 2.44 4.88 8.34 3.51 4.79
CoOp 71.44 63.55 45.76 47.81 73.74 57.72 1.10 4.19 8.40 15.34 0.80 7.18
+DOR(ours) 71.47 64.47 48.28 50.12 76.05 59.73 1.64 1.95 4.97 11.07 1.58 4.89
MaPLe 72.05 64.57 48.78 47.66 76.61 59.41 1.13 2.56 4.88 12.42 1.06 5.23
+DOR(ours) 71.89 64.94 48.77 48.29 76.20 59.55 1.46 1.89 3.96 11.08 1.37 4.58
DOR is insensitive to hyperparameters.
To further illustrate the influence of hyperparameter $\lambda$, we present a sensitivity analysis. We report the average performance on 11 datasets under base-to-new evaluation protocol. As is shown in in Figure 4, we can observe that DOR demonstrates robustness in model calibration as $\lambda$ in Eq.(10) varies. Although KgCoOp has better calibration on new classes as $\lambda$ increases, it sacrifices accuracy and calibration on base classes while our method does not. The results verify that DOR is an effective approach to boosting calibration performance on new classes while maintaining performance on base classes.
DOR is robust to covariate shifts.
To comprehensively verify the robustness of DOR, we further evaluate its performance in the domain generalization setting, i.e., there is a covariate shift between training and testing datasets. Specifically, we first fine-tune the model with all classes of ImageNet on the 16-shot setting and then evaluate it on 4 types of datasets with covariate shifts. As is shown in Table 3, the calibration performance indicates DOR maintains stability in the presence of covariate shifts. Although DOR is not specially designed for the covariate shift, it outperforms the calibration baseline of CoOp and MaPLe, reducing ECE by 2.29% and 0.65% under distribution shifts respectively. Meanwhile, DOR demonstrates superior accuracy in domain generalization and successfully maintains in-distribution performance.
6 Discussion
Table 4: Calibration results of ECE (%) on fine-tuning of visual representation. DOR-V(ision) is effective for better calibration via visual representation regularization.
@@ 6 Discussion
Method Base New Base New Base New Base New Base New Avg.
@@
Table 5: Calibration results of ECE (%) with different outliers . Oracle* denotes new classes used in the test-time.
Method Source Base New HM
CoOp near-OOD 2.68 7.09 4.89
far-OOD 2.95 7.72 5.34
oracle* 3.13 4.34 3.74
MaPLe near-OOD 3.00 4.52 3.76
far-OOD 2.65 4.95 3.80
oracle* 3.15 4.22 3.69
What makes a good regularization for CLIP fine-tuning?
Based on the above observation, we have demonstrated that using the text outlier as the regularization source could be better than the base classes used in prompt tuning. In particular, we use the relevant but non-overlapped outlier with the fine-tuning task (referred to as near-OOD) as the regularization data. Such selection raises a question: why we prefer to choose near-OOD? To address this, we conducted an ablation on different types of outliers. We considered three types of outliers: near-OOD(ours), far-OOD and new classes used at test time. The far-OOD is selected by applying the opposite operation of Eq. 7. Since target new classes are unknown during fine-tuning, we view them as an oracle, which serves as a performance upper bound for base-to-new tasks. We report the average performance on the base-to-new datasets.
We present the results in Table 5. Since we primarily fine-tune the model for a specific downstream task, selecting far-OOD data completely unrelated to the original task may not be optimal for confidence calibration. Moreover, while the oracle can achieve impressive performance on the calibration of new classes, the fixed number of outliers may cause the model to overfit them, resulting in a decline in generalization. Additionally, we do not know which categories will be used at test time during the fine-tuning phase. Therefore, we dynamically use near-OOD as regularization data, which reduces calibration errors on new classes while preserving performance on base classes.
Can the criterion of DOR be extended to visual tuning?

### v4

@@ 4 Method: Dynamic Outlier Regularization
In the previous analysis, we show that textual divergence is the key for CLIP calibration. To preserve the calibration capacity of zero-shot CLIP, our key idea is to regularize the textual divergence of new classes without restricting that of base classes. To this end, we use textual outliers to regularize fine-tuned CLIP. We provide clear evidence to show that the regularization using textual outliers can easily mitigate the overconfidence in the new classes without altering the fine-tuning objective in Appendix B.2.
@@
With this in mind, we construct a set of textual outliers using nouns from the large language lexical database. Specifically, we mainly use WordNet (Miller, 1995) as the database, which is a large English lexical database containing over 150,000 words. We select nouns from WordNet that do not overlap but share higher-level concept relations with the base classes used in the fine-tuning, and then incorporate them into our regularization term for prompt tuning. We demonstrate the effectiveness of using relevant textual outliers in Table 8.
@@
$s_{i}=\frac{1}{n}\sum_{j=1}^{n}\operatorname{sim}\left(\psi\left(\boldsymbol{t}_{o_{i}}\right),\psi\left(\boldsymbol{t}_{c_{j}}\right)\right),$ (6)
where $\boldsymbol{t}_{o_{i}}$ represents the textual tokens of a noun $o_{i}$ using a fixed prompt like “a photo of a [class-name]”, and $i\in\{1,2,\ldots,m\}$. $\psi$ is the text encoder of zero-shot CLIP. Then we get the set of textual outliers using the score ranking.
@@
$\mathcal{L}_{\text{dor}}=1-\frac{1}{B}\sum_{b=1}^{B}\operatorname{sim}\left(\psi\left(\boldsymbol{t}_{o_{b}}^{\prime}\right),\psi\left(\boldsymbol{t}_{o_{b}}\right)\right),$ (8)
where $\operatorname{sim(\cdot)}$ denotes the cosine similarity function and $\boldsymbol{t}_{o_{b}}$ denotes the token of the textual outlier $o_{b}$. Using the regularization, the textual divergence of the fine-tuned CLIP will be regularized to be consistent with the zero-shot CLIP. Different from KgCoOp (Yao et al., 2023), the outlier regularization does not restrict the textual feature deviation of base classes, which is explicitly shown in Figure 9.
@@
$\mathcal{L}_{total}=\mathcal{L}_{\text{ce}}+\lambda\cdot\mathcal{L}_{\text{dor}},$ (9)
@@
Extension to other robust fine-tuning algorithms It is worth noting that our regularization is a general method and can be easily incorporated into existing regularization-based tuning for CLIP, including knowledge-guided fine-tuning (Yao et al., 2023, 2024), multimodal consistency (Khattak et al., 2023b; Roy & Etemad, 2023), Self-Regularization (Khattak et al., 2023b) etc. Given the existing robust fine-tuning objective $\mathcal{L}_{\text{robust}}$, we formalize the fine-tuning objective as:
$\mathcal{L}_{\text{total }}=\mathcal{L}_{\text{robust}}+\lambda\cdot\mathcal{L}_{\text{dor}}.$ (10)
Noticeably, DOR offers several compelling advantages:
  • Easy-to-use: DOR leverages text outliers for regularization, which is readily available and easy to collect.
  • Algorithm-agnostic: DOR can be easily incorporated into existing fine-tuning methods (See Table 1 and 2) or calibration algorithms for CLIP (See Appendix I). Furthermore, our method can be extended to visual fine-tuning methods with image outliers (See Table 7).
  • Fine-tuning-nontoxic: Compared with existing regularization-based methods, DOR does not conflict with the fine-tuning objective and breaks the calibration trade-off. (See Table 3 and Appendix B.2).
@@ 5 Experiments
Table 1: Average calibration across 11 datasets. “+DOR(Ours)” to our method applied to standard tuning methods. ↓ indicates smaller values are better. Calibration error is given by ×$\times 10^{-2}$. “HM” denotes the harmonic mean. Bold numbers are significantly superior results.
@@
Table 2: Average ECE (%) of regularization-based methods across 11 datasets. Bold numbers are significantly superior results. DOR constantly improves the calibration when incorporated with existing regularization-based methods.
Metric KgCoOp TCP PromptSRC CoPrompt PromptKD
Vanilla +DOR Vanilla +DOR Vanilla +DOR Vanilla +DOR Vanilla +DOR
Base 5.82 6.07 4.71 4.79 3.7527 3.88 2.56 2.96 4.73 4.81
New 4.48 3.99 4.07 3.80 4.15 3.80 5.96 4.69 4.38 3.66
HM 5.15 5.03 4.39 4.29 3.95 3.84 4.26 3.83 4.56 4.24
5.1 Experimental Setup
Benchmark setting. Following recent works (Zhou et al., 2022b; Wang et al., 2024), we perform two evaluations in two standard benchmark settings: 1) Generalization from Base-to-New Classes: A downstream dataset will be equally split into base and new classes. The model is trained only on the base classes in a few-shot setting and evaluated on base and new classes. In practice, we mainly focus on the harmonic mean value from both classes. 2) Domain Generalization: The modal is trained on ImageNet-1k in a few-shot manner and evaluated on four other ImageNet datasets that contain various types of domain shifts.
Datasets. For the base-to-new evaluation, we use 11 image recognition datasets that cover diverse classification tasks including ImageNet (Deng et al., 2009), Caltech101 (Fei-Fei et al., 2004), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), FGVCAircraft (Maji et al., 2013), SUN397 (Xiao et al., 2010), UCF101 (Soomro et al., 2012), DTD (Cimpoi et al., 2014) and EuroSAT (Helber et al., 2019). For domain generalization, we use ImageNet-1k as the source dataset and its four variants as target datasets including ImageNetV2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b), and ImageNet-R (Hendrycks et al., 2021a).
Baselines. We compare our method with standard prompt tuning algorithms including CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), MaPLe (Khattak et al., 2023a) and DEPT (Zhang et al., 2024b). We also incorporate DOR with regularization-based tuning including KgCoOp (Yao et al., 2023), TCP (Yao et al., 2024), PromptSRC (Khattak et al., 2023b), CoPrompt (Roy & Etemad, 2023) and PromptKD (Li et al., 2024).
Implementation details. We use CLIP ( ViT-B/16) (Radford et al., 2021) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022a). We list the details of the compared methods in Appendix D.
Evaluation metrics. We use 4 standard metrics of confidence calibration in our evaluation including Expected Calibration Error (ECE) (Guo et al., 2017), Maximum Calibration Error (MCE) (Guo et al., 2017), Adaptive Calibration Error (ACE) (Nixon et al., 2019) and Proximity-Informed Expected Calibration Error (PIECE) (Xiong et al., 2023).
@@ 5.2 Results
Table 3: Average calibration results of ECE (%) using various regularization on 11 datasets. DOR breaks the calibration trade-off between base and novel classes.
Method Variant Base New HM
CoOp Vanilla 3.07 14.49 8.78
+KG 5.82 4.48 5.15
+DOR 2.47 6.48 4.47
MaPLe Vanilla 2.75 5.46 4.11
+KG 4.01 4.29 4.15
+DOR 3.06 4.26 3.66
CoPrompt Vanilla 2.60 5.96 4.28
+KG 4.01 4.99 4.50
+DOR 2.98 5.14 4.06
Table 4: Average calibration results of ECE (%) on 11 datasets. DOR denotes that the outlier pool excludes all new classes.
Method Variant Base New HM
CoOp Vanilla 3.07 14.49 8.78
+DOR 2.67 6.49 4.58
+DOR 2.82 6.77 4.80
MaPLe Vanilla 2.75 5.46 4.11
+DOR 2.83 4.44 3.64
+DOR 2.89 4.51 3.70
CoPrompt Vanilla 2.60 5.96 4.28
+DOR 2.96 4.69 3.83
+DOR 2.71 4.93 3.82
@@
DOR enhances the calibration of existing prompt-tuning methods. Table 1 shows the calibration performance of 4 standard tuning baselines w/ or w/o our DOR. We find that DOR can consistently reduce the calibration error on new classes. For instance, DOR significantly reduces the ECE from 14.58% to 6.49% on new classes and maintains the ECE from 3.07% to 2.67% on base classes, which makes CoOp more reliable in terms of predicted confidence. Moreover, we incorporate DOR with other robust tuning methods in Table 2. We observe a trade-off between the base and new class for these methods, e.g., CoOp outperforms TCP on base classes but significantly underperforms TCP on new classes. Notably, DOR can continually reduce miscalibration on new classes across various metrics without calibration trade-offs on base and new classes. In summary, our proposed DOR can consistently boost calibration performance on new classes upon existing state-of-the-art prompt tuning methods without compromising the vanilla fine-tuning objectives. Due to space constraints, we provide detailed calibration results of all datasets in Appendix F. Moreover, to illustrate how DOR affect the probability distribution, we provide a visualization in Appendix G.
DOR effectively breaks the calibration trade-off. To directly demonstrate the role of DOR in the regularization during CLIP fine-tuning, we compare DOR and KgCoOp on 3 baselines (CoOp, MaPLe, and CoPrompt), and present the results in the table 3. The results demonstrate that integrating with DOR can consistently outperform those with KgCoOp on overall performance. In particular, our method achieves much better performance than KgCoOp on Base classes, while KgCoOp performs well on New classes. This is consistent with the analysis presented in Subsection 3.2: KgCoOp anchors the confidence level, leading to underconfidence on Base classes. In short, the results demonstrate the superiority of DOR in breaking the calibration trade-off between base and novel classes.
DOR do not rely on the overlap with new classes. As we mentioned in Section 4, DOR samples leverages text outliers for regularization. To further analyze the potential overlap with new classes used in the test time, we exclude all new classes from the outlier pool (denoted as DOR). As is shown in Table 4, the average results on 11 datasets show that DOR without overlap achieves comparable performance to DOR, significantly improving the performance on all three baselines. Therefore, the effectiveness of our method does not rely on the overlap with new classes. Moreover, we present all selected outlier texts for 6 datasets, showing almost no class overlap, especially on StanfordCars and FGVCAircraft in Table 10. In summary, leverages general semantic information from language space rather than memorizing target classes, which ensures its fairness and generalizability.
@@
Table 5: Accuracy comparison on domain generalization datasets. DOR boosts the calibration and generalization of existing methods.
ECE (↓) Accuracy (↑)
@@
CLIP 1.86 2.44 4.88 8.34 3.51 4.79 66.73 60.87 46.09 47.81 73.98 57.19
CoOp 1.10 4.19 8.40 15.34 0.80 7.18 71.44 63.55 45.76 47.81 73.74 57.72
+DOR(ours) 1.64 1.95 4.97 11.07 1.58 4.89 71.47 64.47 48.28 50.12 76.05 59.73
MaPLe 1.13 2.56 4.88 12.42 1.06 5.23 72.05 64.57 48.78 47.66 76.61 59.41
+DOR(ours) 1.46 1.89 3.96 11.08 1.37 4.58 71.93 64.94 48.77 48.29 76.20 59.55
Table 6: Average accuracy (%) across 11 base-to-new datasets. “Vanilla” denotes the baseline w/o DOR. DOR can improve the generalization capacity on unseen classes while maintaining the performance on base classes.
ZSCLIP CoOp CoCoOp MaPLe KgCoOp DEPT CoPrompt PromptKD
Class Vanilla Vanilla +DOR Vanilla +DOR Vanilla +DOR Vanilla +DOR Vanilla +DOR Vanilla +DOR Vanilla +DOR
Base 69.49 82.97 83.20 80.57 79.89 82.11 82.08 82.29 82.13 83.70 83.81 82.32 82.39 85.74 85.52
New 74.32 61.74 72.01 72.47 74.59 73.89 75.89 72.21 73.14 65.04 71.39 73.29 74.50 79.80 80.81
HM 71.90 72.36 77.61 76.52 77.24 78.00 78.98 77.25 77.64 74.37 77.60 77.81 78.44 82.77 83.17
DOR benefits base-to-new generalization. To further verify that our DOR is effective on new classes and non-toxic for performance on base classes, we summarize the comparison of average test accuracy in Table 6. Similar to the evaluation of calibration, a salient observation is that our proposed DOR drastically improves base-to-new generalization, with its accuracy consistently outperforming all existing baselines in the harmonic mean of base and new classes. Moreover, DOR almost entirely preserves model capability in the classification performance on base classes without degeneration. For instance, applying DOR in DEPT can increase accuracy on new classes from 65.04% to 71.39%, while keeping the accuracy of base classes similar to the baseline. Given that most prompt tuning methods lag behind zero-shot CLIP on the accuracy of new classes, an intuitive explanation is that DOR aligns the features of unseen classes with the zero-shot features, which can preserve the zero-shot generalization on new classes. In addition, we observe that some methods (e.g., CoCoOp and MaPLe) with DOR, resulting in improvements of 2.12% and 2.00% respectively, outperforming zero-shot accuracy on new classes. This demonstrates that DOR can significantly enhance the base-to-new generalization capacity of fine-tuned CLIP. To further demonstrate DOR can be applied to domain-specific open-vocabulary tasks, we provide more results on medical datasets in Appendix J.
Table 7: Calibration results of ECE (%) on fine-tuning of visual representation. DOR-V(ision) is effective for better calibration via visual representation regularization. HM refers to harmonic mean.
@@
Method Base New Base New Base New Base New Base New HM
@@
Table 8: Calibration results of ECE (%) with different outliers. Oracle* denotes new classes that meet during test time.
w/o Near Far Random Oracle*
Base 3.07 2.68 2.95 2.80 3.13
New 14.58 7.09 7.72 7.33 4.34
HM 8.83 4.89 5.34 5.07 3.74
Figure 4: Parameter sensitivity between KgCoOp (KG) and DOR (w/ CoOp). Compared with KG, the accuracy and ECE of DOR are not sensitive to $\lambda$ base classes. Left: Accuracy. Right: ECE.
DOR is robust to covariate shifts. To comprehensively verify the robustness of DOR, we further evaluate its performance in the domain generalization setting, i.e., there is a covariate shift between training and testing datasets. Specifically, we first fine-tune the model with all classes of ImageNet on the 16-shot setting and then evaluate it on 4 types of datasets with covariate shifts. As shown in Table 5, the calibration performance indicates DOR maintains stability in the presence of covariate shifts. Although DOR is not specially designed for the covariate shift, it outperforms the calibration baseline of CoOp and MaPLe, reducing ECE by 2.29% and 0.65% under distribution shifts respectively. Meanwhile, DOR demonstrates superior accuracy in domain generalization and successfully maintains in-distribution performance. Such regularization will not hinder the generalization capacity of fine-tuned CLIP.
DOR is insensitive to hyperparameters. To further illustrate the influence of hyperparameter $\lambda$, we present a sensitivity analysis. We report the average performance on 11 datasets under the base-to-new evaluation protocol. As shown in Figure 4, we can observe that DOR demonstrates robustness in model calibration as $\lambda$ in Eq.(10) varies. Although KgCoOp has better calibration on new classes as $\lambda$ increases, it sacrifices accuracy and calibration on base classes while our method does not. The results verify that DOR is an effective approach to boosting calibration performance on new classes while maintaining performance on base classes. We provide more ablations on DOR including selection strategy, similarity metric, outlier numbers, update frequency, and outliers databases in Appendix H.
6 Discussion